# FMNV: A Dataset of Media-Published News Videos for Fake News Detection

### v2

@@ Abstract
News media, particularly video-based platforms, have become deeply embedded in daily life, concurrently amplifying risks of misinformation dissemination. Consequently, multimodal fake news detection has garnered significant research attention. However, existing datasets predominantly comprise user-generated videos characterized by crude editing and limited public engagement, whereas professionally crafted fake news videos disseminated by media outlets—often politically or virally motivated—pose substantially greater societal harm. To address this gap, we construct FMNV, a novel dataset exclusively composed of news videos published by media organizations. Through empirical analysis of existing datasets and our curated collection, we categorize fake news videos into four distinct types. Building upon this taxonomy, we employ Large Language Models (LLMs) to automatically generate deceptive content by manipulating authentic media-published news videos. Furthermore, we propose FMNVD, a baseline model featuring a dual-stream architecture integrating CLIP and Faster R-CNN for video feature extraction, enhanced by co-attention mechanisms for feature refinement and multimodal aggregation. Comparative experiments demonstrate both the generalization capability of FMNV across multiple baselines and the superior detection efficacy of FMNVD. This work establishes critical benchmarks for detecting high-impact fake news in media ecosystems while advancing methodologies for cross-modal inconsistency analysis. Our dataset is available in https://github.com/DennisIW/FMNV.
@@ 1 Introduction
With the rapid evolution of social media into video-centric content ecosystems, multimodal fake news has emerged as a novel information threat in the digital age. By seamlessly integrating textual narratives, audiovisual materials, and contextualized editing techniques, such content constructs highly deceptive narrative [1, 2, 3, 4, 5], exhibiting exponentially greater propagation efficacy compared to traditional unimodal forms. Recent studies reveal that 73% of disinformation on social platforms in 2023 employed cross-modal coordinated manipulation strategies [6], posing severe challenges to public cognitive security. Multimodal Fake News Detection (MFND), as a pivotal task in cross-modal intelligent computing, aims to achieve precise identification of deceptive content through in-depth exploration of semantic correlations and contradictory patterns across text, visual, auditory, and other heterogeneous data sources [2, 12].
The rise of social media based on video has gradually shifted the focus of multi-modal false news detection to the detection of false news videos. However, the existing video data sets all contain many poorly edited videos posted by individual users, which attract a small amount of attention and are easy for people to distinguish the truth and falseness of news, so there is little social harm. However, because the news media receives more attention, its fake news videos that are driven by political propaganda or traffic are more harmful. To address this challenge, we propose Fake Media News Videos (FMNV), a dataset consisting of 2,393 news videos released by news media enhanced with large language model data. Specifically, we conduct an empirical analysis to categorize fake news videos and give examples. Then, we collect 2,393 news videos published by news media on YouTube and Twitter. Since the workload of human annotation is too large and the accuracy is not high, and most of the videos published by media are real news videos, there is a problem of data imbalance, so we adopt Large Language Models (LLMs) for data enhancement. 1,500 of the news videos are changed to different categories of fake news videos. Each video contains three types of modal information: title, video clip and audio. Fake news videos have the problem of inconsistent information between different modals or consistent information between modals but contrary to common sense.
We also propose Fake Media News Video Detection (FMNVD) as the baseline for FMNV. This model employs a dual-stream architecture integrating CLIP and Faster R-CNN to capture discriminative video features, followed by a co-attention layer for multimodal aggregation. Numerical experiments conducted on FakeTT and FMNV datasets using existing methods and FMNVD demonstrate the effectiveness of our proposed framework.
The main contributions are summarized as follows:
  • $\bullet$
    We conduct an empirical analysis to categorize fake news videos in video datasets and the real world, giving examples of each category.
  • $\bullet$
    We construct FMNV, a novel fake news video detection dataset that contains only news videos published by news media. And we use LLMs data augmentation instead of manual annotation, solving the problem of data imbalance.
  • $\bullet$
    We propose FMNVD, a baseline model that leverages CLIP and Faster R-CNN for feature extraction and employs co-attention for feature aggregation. Comparative experiments validate the effectiveness of FMNV.
2 Related Work
The field of multimodal fake news detection (MFND) has witnessed significant progress through the development of diverse datasets. Early research primarily focused on image-text multimodal scenarios, leading to several influential datasets. FVC-2018 [8] pioneered multimodal verification by incorporating images, text, and metadata from news articles. FakeNewsNet [9] further enriched the landscape with social context features including user profiles and propagation networks. Subsequent innovations introduced specialized datasets: VAVD [10] addressed visual claim verification through image-claim pairs, while LIAR [11] emphasized textual metadata analysis with speaker-specific features. The recently proposed MCFEND [12] advanced cross-modal fusion capabilities through its multi-granularity framework. These datasets collectively established foundational resources for analyzing textual, visual, and social modalities in static content.
However, the emergence of video-centric social platforms has necessitated a paradigm shift toward dynamic multimodal analysis. Researchers have responded by developing video-specific MFND datasets that account for temporal and audiovisual complexities. A notable contribution comes from Qi et al. [3], who constructed FakeSV - a comprehensive Chinese short video dataset containing 5,538 videos from Douyin and Kuaishou platforms. This dataset uniquely categorizes content into 1,827 real videos, 1,827 fake videos, and 1,884 debunked videos across 738 events. Each entry provides rich multimodal features including video content, titles, metadata (upload dates/locations), user comments, and creator profiles. In 2024, Bu and Qi et al. [14], based on FakeSV, proposed FakeTT, an annotated English video dataset with audiovisual text inconsistencies from TikTok. These video datasets emphasize dynamic multimodal interactions, temporal inconsistency detection, and platform-native engagement features, providing critical infrastructure for detecting evolving deception tactics in video-first media environments.
@@ 3 The FMNV Dataset
In this section, we construct a comprehensive dataset through a multi-stage process, integrating empirical analysis, real-world data collection, large-model-assisted augmentation, and in-depth dataset evaluation. First, we examine the semantic inconsistency patterns commonly found in fake news videos to guide data annotation. Then, we collect and curate video samples from various sources, followed by data augmentation using large language models and shot segmentation. Finally, we perform quantitative and qualitative analysis, comparing our dataset with existing benchmarks to highlight its quality and diversity.
@@ 3.1 Empirical Analysis
Existing communication studies [15, 16] suggest that fake news videos can be classified from multiple dimensions, such as production motives, technical means, and content nature. We analyze both public fake news video datasets and real-world collected samples. Fake news often misleads through edited and spliced footage or emotionally manipulative, contradictory content. Since news videos typically contain three modalities—title, video, and audio—we examine the semantic differences among them. When one modality conflicts with the others or defies common sense, the video is more likely to be fake. Based on modality-level inconsistency, we categorize fake news videos into four types:
@@ 3.1.1 Contextual Dishonesty
Titles, as the video’s summary, significantly shape audience perception. Malicious title manipulation causes semantic dissonance by omitting context or using emotionally charged language. Common tactics include subject-object substitution, exaggeration, and removal of qualifiers—misleading readers who engage only with headlines.
@@ 3.1.2 Cherry-picked Editing
Video content is often altered by omitting key scenes or splicing unrelated footage to construct a false narrative. Tactics include removing moderate remarks, excluding peaceful scenes, or combining unrelated events, thereby amplifying bias and misleading viewers.
@@ 3.1.3 Synthetic Voiceover
These videos pair AI-generated or cloned speech with unrelated visuals. For example, a fake announcement might play over footage of a harmless event. The mismatch exploits viewers’ assumptions of audiovisual coherence, often bypassing detection through professional editing.
@@ 3.1.4 Contrived Absurdity
Such videos use consistent multimodal content to present absurd claims (e.g., “alien bodies found” or “5-ton dog”), often generated by AI or CGI. Despite surface coherence, the narratives defy logic, aiming to attract attention through curiosity and shock.
@@ 3.2 Data Collection
At present, the existing fake news video detection datasets are collected from the short video platforms Douyin and Tiktok. These videos are mainly published by individual users with a small number of followers. Moreover, the clips of fake news videos are poor and people can easily distinguish the authenticity from the fake ones, so this type of fake news videos is not too harmful. The news video released by the news media is well-produced and will receive more attention, and the news media may also publish fake news videos for political purposes or commercial interests, which is particularly harmful. To solve this problem, we choose to collect news videos posted by news media from Twitter and YouTube. Specifically, we use web crawler technology to collect videos on 12 topics including accidents, epidemics and politics published by 27 news media on Twitter and YouTube in the past five years. In order to minimize noise and improve the quality of the dataset, we manually check and remove the videos with blurred images and too short duration. A total of 2,393 news videos are retained. These news videos contain information in three modals: title, video clip and audio, and the semantic information expressed in all three modals is the same.
@@ 3.3 Data Augmentation
Since our news videos are collected from news media, most of the videos are real, so there is a problem of data imbalance. Therefore, inspired by Liu et al. [13], we adopt LLMs to generate fake news videos from original real news videos. According to the four categories of fake news videos in the empirical analysis, we classify the fake news videos in FakeTT. Then we use different methods to generate the corresponding four different categories of fake news videos, and make the proportion of the four types similar to that in FakeTT, resulting in CD:600, CE:450, SV:300, CA:150.
@@
Contextual Dishonesty (CD) Since the semantic information of the title mismatches other modalities, we utilize LLMs to modify the original video titles, ensuring inconsistency with the semantic content of the video. Specifically, we employ ERNIE 4.0 to regenerate video titles with divergent semantics. We design prompts to instruct the LLMs to alter either the object, verb phrase, or generate different events under the same topic in the titles. Table 1 illustrates our prompts and examples of modified titles. Subsequently, we manually review and remove redundant information from the generated responses, refining poorly constructed instances. This process yielded CD-type fake news videos with semantically mismatched titles.
Cherry-picked Editing (CE) Semantic inconsistencies caused by modifications in the video modality primarily arise from deleting key frames of the original material and splicing different segments. To simulate this forgery process, we adopt a similar approach. For selecting critical segments in the video, we use TransNetV2  [21] to perform shot segmentation on the video clips. Subsequently, frames are extracted from each video segment at two-seconds intervals. The video’s title and all extracted frames from each segment are then mapped into a shared vector space using CLIP. The cosine similarity between each frame and the title is calculated, and the average similarity across all frames in a segment is assigned as the segment score. After comparing the scores of all segments, we remove the video segments with the top 1/3 of the segment score. Due to the loss of critical information, these videos exhibit logical incoherence and modality mismatch, thereby being classified as CE-type fake news videos.
@@
Synthetic Voiceover (SV) Semantic contradictions in audio modals are artificially created by generating audio content that diverges from the video’s original topic. To achieve this, we leverage LLMs to synthesize textually unrelated narratives under the same topic. Specifically, the video’s title and image captions of key frames are fed into ERNIE with structured prompts to generate topic-coherent but semantically irrelevant news articles. These generated texts are then converted into speech using text-to-speech tools VITS to replace the original audio. The resulting videos exhibit cross-modal contradictions between audio narration and visual content, classified as SV-type fake news videos.
Contrived Absurdity (CA) This type of fake news video maintains semantic consistency across modalities but violates common sense through illogical or absurd claims. Adopting a method similar to the CD-type processing, we utilize ERNIE 4.0 to exaggerate the video title in an illogical way while preserving their original meaning. The prompts instruct the LLMs to amplify specific elements (e.g., inflating quantities or distorting causality) without altering core entities or events. In this way, we get CA-type fake news videos with the same semantic information but contrary to common sense.
@@
Following data augmentation, we end up with a final dataset comprising five categories of 2393 videos.
@@ 3.4 Data Analysis
The FMNV dataset comprises a total of 2,393 news videos, including 893 videos categorized as real and 1,500 videos categorized as fake. The fake category can be further divided into four subtypes: CD: 600 videos with title modifications causing semantic inconsistencies; CE: 450 videos where key video clips matching the title were deleted; SV: 300 videos with audio replaced by reports of different events under the same theme; CA: 150 videos featuring titles with synonymous but exaggerated narration.
When compared to other MFND video datasets, FMNV offers notable advantages. In our LLM text generation process for data augmentation, we use a word substitution strategy to replace some words from the original text to change the original meaning. This ensures that the length of both positive and negative texts remains comparable, thereby preventing significant length disparities. With an average video duration of 73.8 seconds, FMNV boasts considerably longer videos than other datasets, and nearly half of the videos are long videos longer than 80 seconds, allowing for a richer information content per video. A comparative analysis of MFND datasets is presented in Table 2.
@@ 4 Method
Fig. 5 illustrates the overall architecture of the proposed FMNVD. Our baseline model consists of two core components: feature extraction and feature aggregation. The feature extraction module utilizes BERT to process textual features from video titles and audio transcripts, while employing a dual-stream architecture with Faster R-CNN and CLIP to capture both local object features and global visual representations from video frames. The feature aggregation module integrates these multimodal features through co-attention mechanisms that model cross-modal correlations, ultimately fusing the enhanced representations for fake news video classification.
@@ 4.1.1 Title
In the title feature extraction module, the video title text is first tokenized using the BERT-base WordPiece tokenizer with added [CLS] and [SEP] tokens. The token sequence is then fed into a 12-layer Transformer encoder with positional embeddings. We extract the final hidden states of all tokens from the last Transformer layer, where the [CLS] token’s 768-dimensional embedding serves as the global semantic representation. To handle variable-length inputs, sequences are padded/truncated to a fixed length $l_{t}=32$. The resulting text features $F_{t}=[w_{1},…,w_{l_{t}}]$ capture contextualized word-level semantics for downstream fusion.
@@ 4.1.2 Video
We employ temporal sampling to extract key frames from video sequences. Specifically, given a video $V$ with duration $T$, we uniformly sample $N$ frames at intervals of $\Delta t$ seconds. Each frame $I_{t}$ is resized to $H\times W$ resolution with RGB normalization.
Our dual-stream architecture synergistically combines local object details and global contextual patterns. For the local object features, Faster R-CNN with ResNet-101-FPN backbone processes each frame to generate region proposals through Region Proposal Network (RPN). We retain the top-$K$ proposals with confidence scores exceeding $\tau=0.5$, with $K=50$. followed by RoI pooling to extract fixed-size features $\{o_{i}\}_{i=1}^{K}\in{R}^{256}$. For global context features, CLIP (ViT-B/32) encoder partitions each frame into a $3\times 3$ grid of patches, yielding 9 patch embeddings $\{p_{j}\}_{j=1}^{9}\in{R}^{512}$ through linear projection.
Before fusion, we perform cross-modal feature alignment: 1) Project local features to CLIP’s semantic space via learnable matrix $W_{a}\in{R}^{512\times 256}$, 2) Apply element-wise L2-normalization to both modalities:
$\hat{o}_{i}=\frac{W_{a}o_{i}}{\|W_{a}o_{i}\|_{2}},\quad\hat{p}_{j}=\frac{p_{j}}{\|p_{j}\|_{2}}$ (1)
The final fused representation concatenates aligned features followed by dimension compression:
$F_{f}=\phi\left(\left[\hat{o}_{1},\dots,\hat{o}_{K}\oplus\hat{p}_{1},\dots,\hat{p}_{9}\right]\right)\in{R}^{d}$ (2)
where $\oplus$ denotes concatenation, $\phi:{R}^{(K+9)\times 512}\to{R}^{d}$ is a trainable projection layer (implemented as two-layer MLP with ReLU), and $d$ is the unified frames feature dimension.
To aggregate frame-wise features into a unified video representation, we employ a two-stage spatial-temporal fusion strategy. First, all frame features are concatenated along the temporal dimension to preserve raw forensic patterns. Then, a learnable compression module projects the high-dimensional features into a compact embedding space.
The video representation is formally obtained through:
$F_{v}=\psi\left(\bigoplus_{t=1}^{N}F_{f}^{(t)}\right)$ (3)
where $\bigoplus$ denotes full concatenation of $N$ frame features, and $\psi:{R}^{N\times 512}\to{R}^{d}$ represents the dimension reduction network implemented as:
$\psi(X)=W_{2}(\text{ReLU}(W_{1}X+b_{1}))+b_{2}$ (4)
with $W_{1}\in{R}^{k\times 512N}$, $W_{2}\in{R}^{d\times k}$ forming a bottleneck structure. Layer normalization is applied before each linear transformation to stabilize training.

### v3

@@ Abstract
News media, particularly video-based platforms, have become deeply embed-ded in daily life, concurrently amplifying the risks of misinformation dissem-ination. Consequently, multimodal fake news detection has garnered signifi-cant research attention. However, existing datasets predominantly comprise user-generated videos characterized by crude editing and limited public en-gagement, whereas professionally crafted fake news videos disseminated by media outlets—often politically or virally motivated—pose substantially greater societal harm. To address this gap, we construct FMNV, a novel da-taset exclusively composed of news videos published by media organizations. Through empirical analysis of existing datasets and our curated collection, we categorize fake news videos into four distinct types. Building upon this taxonomy, we employ Large Language Models (LLMs) to automatically generate deceptive content by manipulating authentic media-published news videos. Furthermore, we propose FMNVD, a baseline model featuring a dual-stream architecture that integrates spatio-temporal motion features from a 3D ResNeXt-101 backbone and static visual semantics from CLIP. The two streams are fused via an attention-based mechanism, while co-attention modules refine the visual, textual, and audio features for effective multi-modal aggregation. Comparative experiments demonstrate both the generali-zation capability of FMNV across multiple baselines and the superior detec-tion efficacy of FMNVD. This work establishes critical benchmarks for de-tecting high-impact fake news in media ecosystems while advancing meth-odologies for cross-modal inconsistency analysis. Our dataset is available in https://github.com/DennisIW/FMNV.
@@ 1 Introduction
The proliferation of video-centric social media platforms has transformed the in-formation landscape, giving rise to increasingly sophisticated forms of multimodal disinformation. Unlike traditional text- or image-based fake news, these video-based narratives integrate textual, visual, and auditory elements to create highly persuasive content [1, 2]. Such content exhibits stronger emotional appeal, greater virality, and heightened resistance to detection, making it a growing threat to public opinion and social trust. Recent studies indicate that over 70% of disinformation incidents on social platforms in 2023 involved coordinated cross-modal manipulation [6], underscoring the urgent need for robust multimodal fake news detection (MFND) systems capable of handling complex and dynamic video data.
Prior work in MFND has primarily focused on static multimodal content, such as image-text combinations, resulting in several influential datasets including FVC-2018 [8], FakeNewsNet [9], and MCFEND [7]. More recent efforts have shifted toward video-based datasets to accommodate the evolving nature of disinformation. Notable among them is FakeSV [3], a Chinese short video dataset comprising over 5,000 labeled videos across real, fake, and debunked categories. Building on this, FakeTT [13] extended the research to English-language platforms like TikTok, focusing on multimodal inconsistencies. These resources have significantly advanced the field by introducing temporal dynamics and audiovisual complexity into the analysis.
However, these datasets mainly consist of user-generated content, which often lacks wide influence and is easier to verify. In contrast, fake news from mainstream media—driven by political or economic agendas—is more polished, widespread, and impactful, yet remains underrepresented. Additionally, large-scale manual annotation is costly and prone to error, and existing datasets suffer from class imbalance skewed toward real news. To address these issues, we introduce Fake Media News Videos (FMNV), a dataset of 2,393 professionally produced news videos sourced from verified YouTube and Twitter accounts. To mitigate annotation and imbalance challenges, we employ large language models (LLMs) to generate 1,500 synthetic fake news videos based on empirically defined misinformation categories. Each video features multimodal content—titles, audio, and visual clips—with cross-modal inconsistencies or violations of common sense.
We also propose Fake Media News Video Detection (FMNVD), a dual-stream baseline model that integrates CLIP and 3D ResNeXt-101 for visual feature extrac-tion, with attention-based gated fusion and co-attention for cross-modal alignment. Experiments on FMNV validate the effectiveness of our framework in detecting complex multimodal disinformation.
@@ 2 The FMNV Dataset
In this section, we construct a comprehensive dataset through empirical analysis, real-world data collection, LLM-assisted augmentation, and dataset evaluation. We first analyze common semantic inconsistency patterns in fake news videos to guide annotation. Then, we collect and curate video samples from various sources, aug-ment them using large language models and shot segmentation, and finally conduct quantitative and qualitative analysis to compare our dataset with existing bench-marks in terms of quality and diversity.
@@ 2.1 Empirical Analysis
Existing communication studies [14, 15] suggest that fake news videos can be classi-fied across multiple dimensions, such as production motives, technical means, and content nature. We analyze both public fake news video datasets and real-world samples. Fake news often misleads through edited footage or emotionally manipula-tive, contradictory content. Since news videos typically contain three modalities—title, video, and audio—we examine the semantic differences among them. When one modality conflicts with the others or defies common sense, the video is more likely to be fake. Based on modality-level inconsistency, we categorize fake news videos into four types:
@@ 2.1.1 Contextual Dishonesty
Titles, as the video’s summary, significantly shape audience perception. Malicious title manipulation causes semantic dissonance by omitting context or using emo-tionally charged language. Common tactics include subject-object substitution, ex-aggeration, and removal of qualifiers—misleading readers who engage only with headlines.
@@ 2.1.2 Cherry-picked Editing
Video content is often altered by omitting key scenes or splicing unrelated footage to construct a false narrative. Tactics include removing moderate remarks, excluding peaceful scenes, or combining unrelated events, thereby amplifying bias and mis-leading viewers.
@@ 2.1.3 Synthetic Voiceover
These videos pair AI-generated or cloned speech with unrelated visuals. For exam-ple, a fake announcement might play over footage of a harmless event. The mis-match exploits viewers’ assumptions of audiovisual coherence, often bypassing de-tection through professional editing.
@@ 2.1.4 Contrived Absurdity
Such videos use consistent multimodal content to present absurd claims (e.g., “alien bodies found” or “5-ton dog”), often generated by AI or CGI. Despite surface coher-ence, the narratives defy logic, aiming to attract attention through curiosity and shock.
@@ 2.2 Data Collection
At present, the existing fake news video detection datasets are collected from the short video platforms Douyin and Tiktok. These videos are mainly published by individual users with a small number of followers. Moreover, the clips of fake news videos are poor and people can easily distinguish the authenticity from the fake ones, so this type of fake news videos is not too harmful. The news video released by the news media is well-produced and will receive more attention, and the news me-dia may also publish fake news videos for political purposes or commercial inter-ests, which is particularly harmful. To solve this problem, we choose to collect news videos posted by news media from Twitter and YouTube. Specifically, we use web crawler technology to collect videos on 12 topics including accidents, epidemics and politics published by 27 news media on Twitter and YouTube in the past five years. In order to minimize noise and improve the quality of the dataset, we manually check and remove the videos with blurred images and too short duration. A total of 2,393 news videos are retained. These news videos contain information in three modals: title, video clip and audio, and the semantic information expressed in all three modals is the same.
@@ 2.3 Data Augmentation
As most collected news videos are from news media, they are predominantly real, resulting in data imbalance. Inspired by Liu et al. [12], we employ LLMs to generate fake news videos based on real ones. Referring to the four categories of fake news videos in our empirical analysis, we generate corresponding types and ensure similar proportions to those in FakeTT: CD:600, CE:450, SV:300, CA:150.
@@
Contextual Dishonesty (CD) To introduce semantic mismatches between titles and other modalities, we employ ERNIE 4.0 to rewrite original video titles, ensuring intentional divergence from the actual video content. Prompts are carefully designed to modify specific elements in the title, such as the object, verb phrase, or to generate alternative events under the same general topic. Examples of prompts and corresponding rewritten titles are shown in Table 1. The generated titles are then manually reviewed to remove redun-dancy and refine poorly structured outputs. This process results in CD-type fake news videos, where the title misrepresents the semantic content of the video.
Cherry-picked Editing (CE) This type introduces semantic inconsistencies through deliberate visual editing. Specifically, we first apply TransNet V2 [17] to perform shot segmentation on the original video, dividing it into a series of coherent visual units. From these seg-ments, a portion is evenly removed and the remaining segments are reassembled to create a shortened version of the video. This simulates a cherry-picking process, where critical or context-defining content is selectively omitted. Although the re-sulting video maintains surface-level visual continuity, it loses essential infor-mation, leading to logical incoherence and misalignment with the original title or audio. These manipulated videos are thus categorized as CE-type fake news, charac-terized by cross-modal inconsistencies arising from selective deletion within the visual stream.
@@
Synthetic Voiceover (SV) Cross-modal contradictions are created in the audio stream by replacing the original narration with unrelated synthesized speech. To achieve this, we use ERNIE 4.0 to generate semantically divergent text while staying loosely on the same topic. The inputs include the video title and image captions of key frames, organized into struc-tured prompts. The generated text is converted to speech using the TTS system VITS, replacing the original audio track. The resulting SV-type fake news videos exhibit clear audio-visual contradictions.
Contrived Absurdity (CA) This type maintains semantic alignment across modalities but introduces absurdity by violating common sense through exaggerated or illogical claims. Following a similar strategy as in CD, we prompt ERNIE 4.0 to rewrite the title in an exaggerated manner that preserves the original meaning but distorts quantities, causality, or logi-cal reasoning. The prompts avoid altering key entities or events, ensuring the seman-tic core remains intact. The final CA-type videos appear internally consistent but convey implausible or nonsensical information.
@@
Following data augmentation, we end up with a final dataset comprising five cat-egories of 2393 videos.
@@ 2.4 Data Analysis
The FMNV dataset comprises a total of 2,393 news videos, including 893 videos categorized as real and 1,500 videos categorized as fake. The fake category can be further divided into four subtypes: CD: 600 videos with title modifications causing semantic inconsistencies; CE: 450 videos where key video clips matching the title were deleted; SV: 300 videos with audio replaced by reports of different events un-der the same theme; CA: 150 videos featuring titles with synonymous but exaggerat-ed narration. When compared to other MFND video datasets, FMNV offers notable advantages. In our LLM text generation process for data augmentation, we use a word substitu-tion strategy to replace some words from the original text to change the original meaning. This ensures that the length of both positive and negative texts remains comparable, thereby preventing significant length disparities. With an average video duration of 73.8 seconds, FMNV boasts considerably longer videos than other da-tasets, and nearly half of the videos are long videos longer than 80 seconds, allowing for a richer information content per video. A comparative analysis of MFND da-tasets is presented in Table 2.
@@ 3 Method
Fig. 5 illustrates the overall architecture of our proposed FMNVD model, com-prising two main components: feature extraction and feature aggregation. We use BERT to encode video titles and audio transcripts, and adopt a dual-stream framework combining 3D ResNeXt-101 and CLIP for dynamic and static visual fea-ture extraction. A co-attention-based aggregation module is then used to integrate multimodal features and perform fake news classification.
@@ 3.1.1 Title
The video title is tokenized using the BERT-base WordPiece tokenizer with [CLS] and [SEP] tokens. The sequence is input to a 12-layer Transformer, and we extract the final hidden state of the [CLS] token as the title representation. To standardize input length, sequences are padded or truncated to $l_{t}=32$. The resulting token embeddings $F_{t}=[w_{1},\dots,w_{l_{t}}]$ serve as contextualized text features.
@@ 3.1.2 Video
We design a dual-stream feature extraction module to jointly capture temporal motion dynamics and static spatial semantics from video sequences. Given an input video $V$ of total duration $T$, we segment it into overlapping clips of 8 frames to extract motion features and sample individual keyframes for static analysis.
For motion representation, we utilize a pre-trained ResNeXt-101 model with 3D convolutions to encode spatio-temporal patterns from short clips. Let $V_{t:t+7}$ denote a segment of 8 consecutive frames; the motion feature is then defined as:
$F_{m}=\psi_{m}V_{t:t+7}\in R^{2048}$ (1)
where $\psi_{m}$ represents the 3D convolutional motion encoder.
In parallel, we extract static visual semantics using the CLIP visual encoder (ViT-B/32), which processes individual sampled frames. The resulting frame-level feature is expressed as:
$F_{s}=\psi_{s}V_{t}\in R^{512}$ (2)
where $\psi_{s}$ denotes the CLIP visual encoder.
To enable fine-grained fusion of dynamic and static information, we introduce an attention-based gated fusion mechanism. Specifically, the motion and static features are linearly projected and concatenated, and a learnable sigmoid gate is applied to compute a dynamic weighting:
$\alpha=\sigma\left(W[F_{m};F_{s}]+b\right)$ (3)
$F_{v}=\alpha F_{m}+(1-\alpha)F_{s}$ (4)
Here, $W$ and $b$ are trainable parameters and $\sigma$ denotes the sigmoid function. The resulting fused visual embedding $F_{v}\in R^{d}$ encodes both short-term dynamics and long-range semantics.